# Dynamic Mixture of Progressive Parameter-Efficient Expert Library for Lifelong Robot Learning

### v1

@@ Abstract
A generalist agent must continuously learn and adapt throughout its lifetime, achieving efficient forward transfer while minimizing catastrophic forgetting. Previous work within the dominant pretrain-then-finetune paradigm has explored parameter-efficient fine-tuning for single-task adaptation, effectively steering a frozen pretrained model with a small number of parameters. However, in the context of lifelong learning, these methods rely on the impractical assumption of a test-time task identifier and restrict knowledge sharing among isolated adapters. To address these limitations, we propose Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) for lifelong robot learning. DMPEL progressively learn a low-rank expert library and employs a lightweight router to dynamically combine experts into an end-to-end policy, facilitating flexible behavior during lifelong adaptation. Moreover, by leveraging the modular structure of the fine-tuned parameters, we introduce coefficient replay to guide the router in accurately retrieving frozen experts for previously encountered tasks, thereby mitigating catastrophic forgetting. This method is significantly more storage- and computationally-efficient than applying demonstration replay to the entire policy. Extensive experiments on the lifelong manipulation benchmark LIBERO demonstrate that our framework outperforms state-of-the-art lifelong learning methods in success rates across continual adaptation, while utilizing minimal trainable parameters and storage.
@@ 1 Introduction
A generalist agent should have the capability to learn and adapt continuously throughout its lifetime, usually termed as lifelong learning [1, 2, 3, 4, 5, 6]. The longstanding challenges are enabling forward transfer (i.e., leveraging knowledge from previous tasks to quickly adapt to new ones) and avoiding catastrophic forgetting (i.e., retaining previously acquired knowledge when learning new tasks) under limited computational and memory capacity.
Due to the notorious sample inefficiency of the traditional tabula rasa approach in robotics, researchers have recently explored the pretrain-then-finetune paradigm [7, 8, 9, 10, 11, 12] that originates from the vision and language domains [13, 14, 15, 16, 17, 18, 19, 20]. Specifically, this paradigm first pretrain a policy, often referred to as a vision-language-action (VLA) model, on large-scale datasets, and then fine-tune it for various downstream tasks. However, in the context of lifelong learning, naively applying full fine-tuning to sequentially arriving robotic tasks can result in severe forgetting and suboptimal performance, failing to meet our requirements. As depicted in Figure 1, prior work in lifelong robot learning can be broadly classified into three distinct approaches. Replay methods [8, 21, 22, 23, 24, 25] retain previous data (e.g., vision-language web data, robot pretraining data, and data from seen tasks) and mix it with in-domain data during training on new tasks. These methods leads to generalizable policies, but also necessitates huge storage space and computational overhead for retaining old samples. Regularization methods [26, 27, 28] balance between new and old tasks by restricting the update of parameters. For example, EWC [27] require an entire copy of frozen model from the previous task and entail substantial computation in estimating the importance of parameters. In addition, they usually struggle in long task sequence due to plasticity-stability dilemma.
In contrast, architectural methods [29, 30, 31] address catastrophic forgetting by explicitly learning task-specific parameters. The widely adopted parameter-efficient fine-tuning (PEFT) techniques [32, 33], such as adapters [34, 35], low-rank adaptation (LoRA) [36], and prompt tuning [37, 38, 39], have proven highly effective for steering frozen foundation models by integrating small, learnable modules in single-task adaptation. In the lifelong learning setting, a practical solution, as exemplified by TAIL [40], involves initializing and training task-specific adapters for each task, subsequently freezing them, and retrieving the appropriate adapter based on the oracle task index. Such method allows efficient adaptation with only a small amount of task-specific parameters, but also exhibits certain limitations. First, assuming test-time task identity is known and retrieving adapter in a hard-coded way could be impractical in real scenarios. Second, failing to effectively share knowledge across isolated adapters result in poor forward transfer. Some prior arts construct a shared adapter pool and enable automatic timestep-level retrieval by, for example, performing query-key matching to select the one with highest similarity from a pool of adapters [41], or using multifaceted prototypes to represent the cluster center in the state space for each adapter [42], but usually suffer from inaccurate retrieval and suboptimal performance.
@@
Figure 1: Existing lifelong learning methods: (a) Replay methods; (b) Regularization methods; (c) Architectural methods; (d) Performance on LIBERO-Object. TAIL with LoRA utilizes significantly fewer resources than ER/EWC with FFT and exhibits no forgetting when provided with a task identifier, but demonstrates lower forward transfer. In contrast, DMPEL leverages a low-rank expert library and coefficient replay (CR$=$5%) to achieve better forward transfer and near-zero forgetting.
@@
In this paper, we introduce Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) to overcome these challenges in lifelong robot learning. DMPEL incrementally builds a library of low-rank experts and uses a lightweight router to dynamically assemble these experts into a unified policy, enabling the agent to adapt flexibly throughout continual adaptation. By exploiting the modularity of fine-tuned parameters, DMPEL incorporates coefficient replay at the router, enabling the policy to accurately integrate experts for all previously seen tasks and thereby effectively preventing catastrophic forgetting. This approach is significantly more efficient in terms of storage and computation than replaying expert demonstrations for the entire policy. Through extensive evaluations on the LIBERO lifelong manipulation benchmark [43], we show that our method achieves higher forward transfer and lower catastrophic forgetting than existing baselines while requiring minimal trainable parameters and storage. Code is publicly released for reproducibility.
@@ 2 Related Work
The traditional tabula rasa paradigm in robotics is widely criticized for its sample inefficiency, prompting researchers to explore the use of foundation models to improve downstream task transfer [44, 45]. One prominent direction involves transferring general visual representations: some studies directly employ foundation models from other domains, such as CLIP [15], SAM [18], and DINOv2 [14], as vision encoders [46, 47], while others pretrain visual representations using self-supervised learning [16, 17] on egocentric video datasets like R3M [48], MVP [49], and VC-1 [50]. Another emerging line of research focuses on pretraining large-scale policies on diverse robotic tasks to capture generalizable behavior priors, for example, RT-1 [7], RT-2 [8], Gato [9], Octo [10], OpenVLA [11], and $\pi_{0}$ [12]. However, when adapting these models to downstream tasks, most approaches either rely on the frozen pretrained model (i.e., zero-shot transfer) or require costly full fine-tuning. Some prior work has explored parameter-efficient fine-tuning (PEFT) by integrating task-specific modules into the pretrained vision encoder [51, 52, 53] or temporal transformer [54, 55, 56]. Despite these advancements, the majority of these methods focus on single-stage adaptation, whereas our approach targets lifelong adaptation to a sequence of new tasks.
@@
Compared to full fine-tuning, PEFT optimizes only a small amount of inserted parameters and keeps the rest untouched, greatly reducing the storage cost of tuned parameters and mitigating the risk of overfitting in low-data regimes [32]. Notably, PEFT techniques share conceptual similarities with architectural methods in lifelong learning. Most studies on lifelong learning with PEFT predominantly utilize prompt-based tuning [37, 39] to adapt frozen pretrained model for tasks such as image classification (e.g., L2P [57], DualPrompt [58], CODA-Prompt [59], and DAP [60]) and text classification (e.g., EPI [61], and Progressive Prompts [62]). Current research primarily investigates strategies for dynamically selecting or generating instance-specific prompts [63]. Recognizing the representational limitations of prompt tuning, subsequent studies have extended to more powerful PEFT techniques, including LAE [64], HiDe-PET [65]. Despite the demonstrated success of PEFT-based lifelong learning in classification tasks, its application remains largely unexplored in robotics. A few pioneering works employ LoRA [36] to store task-specific knowledge, yet diverge in their retrieval mechanisms: TAIL [40] relies on oracle task identifier, L2M [41] implements query-key matching between the current context (the state embedding from the frozen pretrained model) and learnable keys, while IsCiL [42] establishes multifaceted prototypes through K-means clustering in the state space, with input states retrieving the closest LoRA parameters based on proximity. Our method incrementally constructs a parameter-efficient expert library and learns a lightweight router to dynamically retrieve and combine these experts.
@@
The Mixture of Experts (MoE) approach, which involves training multiple specialized models for specific subtasks, has gained significant attention [66, 67, 68]. A key method for combining these experts is parameter ensemble, supported by the linear mode connectivity phenomenon, which shows models converging to a single low-loss basin [69]. Task Arithmetic [70] introduces task vectors, representing the difference between fine-tuned and pretrained models, enhancing multi-task model performance when merged. Follow-up work [71, 72] improves this by learning input-conditioned layer-wise fusion coefficients, increasing flexibility and performance. Model fusion techniques also apply to PEFT, where previously learned adapters are reused for efficient task adaptation [73, 74, 75, 76], including LoRAHub [77] and AdapterFusion [78]. In image generation, merging separate LoRAs facilitates joint style and content generation [79, 80, 81]. In robotics, models like MELA [82], MoE-Loco [83], MORE [84], and SDP [85] create versatile, adaptive behaviors for multiple tasks by fusing expert sets. Our method aims to enhance forward transfer in lifelong robot learning by flexibly fusing previously learned low-rank experts for each sub-module in the policy. Besides, we exploit the modularized design by replaying coefficient on the router to mitigate catastrophic forgetting.
@@ 3.1 Lifelong Robot Learning
Due to the significant challenge of sparse reward in robot manipulation, we consider a more practical lifelong imitation learning setting [43, 40], where a robot sequentially learns over a stream of tasks $\{\mathcal{T}^{1},\cdots,\mathcal{T}^{K}\}$ given an expert demonstration dataset $\mathcal{D}^{k}=\{\tau_{n}^{k}\}_{n=1}^{N}$ for each task $\mathcal{T}^{k}=(d_{0}^{k},l^{k})$, $1\leq k\leq K$. Each expert trajectory $\tau_{n}^{k}$ can be represented as $(o_{0},a_{0},\cdots,o_{\mathcal{H}})$, in which observation $o_{t}$ at each step $t$ includes RGB images and proprioceptive states. Following common pratice in partially observable MDPs, we approximate the current state $s_{t}$ by stacking prior observations, i.e., $s_{t}=o_{\leq t}=(o_{0},\cdots,o_{t})$. We perform behavior cloning [86] to learn a stochastic policy $\pi_{\theta}$ by minimizing the negative log-likelihood loss $L$:
@@ 3.2 Evaluation Metrics
We use three metrics to evaluate the performance of lifelong robot learning [43, 87]: forward transfer (FWT), negative backward transfer (NBT), area under the success rate curve (AUC). Formally, we evaluate the policy on current task $\mathcal{T}^{k}$ when reaching a checkpoint $c\in\mathcal{C}$ (i.e., every multiple epochs of training) and we denote the success rate during evaluation as $s_{k,k,c}$. When the training on current task is completed, we find the earliest checkpoint $c_{k}^{*}$ that achieves the best performance and mark it as the final success rate $s_{k,k}$, i.e., $s_{k,k}=s_{k,k,c_{k}^{*}}=\max_{c\in\mathcal{C}}s_{k,k,c}$. We only keep the best checkpoint for future training and evaluation as if the agent stops learning after $c_{k}^{*}$, i.e., for all $c\geq c_{k}^{*}$, $s_{k,k,c}=s_{k,k}$. This checkpoint $c_{k}^{*}$ will be further evaluated on previous tasks $\mathcal{T}^{j},j\leq k$, where the success rate is denoted as $s_{k,j}$. Then we can define the three metrics as:
@@
Intuitively, higher FWT indicates faster adaptation to new tasks, lower NBT indicates less catastrophic forgetting on old tasks, and higher AUC demonstrates an overall better performance across tasks and a good balance between FWT and NBT.
@@ 4.1 Base Policy Architecture
We use an autoregressive policy [11, 88] similar to the one used in the LIBERO benchmark [40, 43] and focus on developing better lifelong learning algorithms. The policy consists of vision, text, and state encoders, an input modality fusion module, a temporal transformer, and an action head.
@@
The policy input is the historical observations $(o_{t-T},\cdots,o_{t})$ and the language instruction $l=\{l_{i}\}_{i=1}^{L}$, where $T$ is the context length and $L$ is the sentence length. The observation at each step $o_{t}=(I^{1}_{t},\cdots,I^{N_{c}}_{t},p_{t})$ includes RGB images $I^{n_{c}}_{t},1\leq n_{c}\leq N_{c}$ from $N_{c}$ different cameras and low-dimensional proprioceptive states $p_{t}$. In experiments, we use images from two cameras, the agent-view image and the eye-in-hand image. We employ the pretrained CLIP ViT-B/16 model [15] to serve as the vision encoder $\mathcal{E}_{I}$ and the text encoder $\mathcal{E}_{L}$. For the joint states and gripper states, we learn separate linear projection layers $\mathcal{E}_{P}$ to project the low-dimensional states to proprioceptive embeddings. In summary, at each timestep, we obtain image embeddings $f^{v}_{t}$, text embeddings $f^{l}_{t}$, and proprioceptive embeddings $f^{s}_{t}$.
@@
We utilize a feature-wise linear modulation (FiLM) [89] to fuse language embeddings with the image and state embeddings. Generally, given the original feature $x$, the conditional input $z$, and the fusion network $f_{\text{FiLM}}$, we can obtain the modulated feature $x^{\prime}=\gamma\odot x+\beta$ where $\gamma,\beta=f_{\text{FiLM}}(z)$. In our case, the conditional input $z$ refers to the language embedding $f^{l}_{t}$, $\gamma$ and $\beta$ are scaling and shifting vectors having the same size as $x$, and $\odot$ refers to element-wise multiplication.
@@
The causal temporal transformer backbone processes a sequence of multi-modal embeddings to produce a latent vector $g_{t}$ at each decision-making timestep. We employ a Gaussian Mixture Model (GMM)-based output head to compute the multi-modal distribution of manipulation actions [40, 43]. During training, we optimize the negative log-likelihood loss between the action distribution and the ground truth action. At inference time, the robot executes the policy by using the mean of the Gaussian distribution with the highest density for end effectors.
Figure 2: Overview of the proposed method DMPEL.

### v2

@@ Abstract
A generalist agent must continuously learn and adapt throughout its lifetime, achieving efficient forward transfer while minimizing catastrophic forgetting. Previous work within the dominant pretrain-then-finetune paradigm has explored parameter-efficient fine-tuning for single-task adaptation, effectively steering a frozen pretrained model with a small number of parameters. However, in the context of lifelong learning, these methods rely on the impractical assumption of a test-time task identifier and restrict knowledge sharing among isolated adapters. To address these limitations, we propose Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) for lifelong robot learning. DMPEL progressively builds a low-rank expert library and employs a lightweight router to dynamically combine experts into an end-to-end policy, enabling flexible and efficient lifelong forward transfer. Furthermore, by leveraging the modular structure of the fine-tuned parameters, we introduce expert coefficient replay, which guides the router to accurately retrieve frozen experts for previously encountered tasks. This technique mitigates forgetting while being significantly more storage- and computation-efficient than experience replay over the entire policy. Extensive experiments on the lifelong robot learning benchmark LIBERO demonstrate that our framework outperforms state-of-the-art lifelong learning methods in success rates during continual adaptation, while utilizing minimal trainable parameters and storage. Code is publicly released for reproducibility.
@@ 1 Introduction
A generalist agent should have the capability to learn and adapt continuously throughout its lifetime, usually termed as lifelong learning (De Lange et al., 2021; Masana et al., 2022; Mendez & Eaton, 2023; Wang et al., 2024a). The longstanding challenges are enabling forward transfer (i.e., leveraging knowledge from previous tasks to quickly adapt to new ones) and avoiding catastrophic forgetting (i.e., retaining previously acquired knowledge when learning new tasks) under limited computational and memory capacity.
Due to the notorious sample inefficiency of the traditional tabula rasa approach in robotics, researchers have recently explored the pretrain-then-finetune paradigm (Brohan et al., 2022; 2023; Kim et al., 2025; Black et al., 2024) that originates from the vision and language domains (Bommasani et al., 2021; Oquab et al., 2024; Radford et al., 2021; Brown et al., 2020). Specifically, this paradigm first pretrain a policy, often referred to as a vision-language-action (VLA) model, on large-scale datasets, and then fine-tune it for various downstream tasks.
@@
Figure 1: Existing lifelong learning methods: (a) Replay methods; (b) Regularization methods; (c) Architectural methods; (d) Performance on LIBERO-Object. TAIL with LoRA utilizes significantly fewer resources than ER/EWC with FFT and exhibits no forgetting when provided with a task identifier, but demonstrates lower forward transfer. In contrast, DMPEL leverages a low-rank expert library and coefficient replay (CR==5%) to achieve better forward transfer and near-zero forgetting.
However, in the context of lifelong learning, naively applying full fine-tuning to sequentially arriving robotic tasks can result in severe forgetting and suboptimal performance, failing to meet our requirements. As depicted in Figure 1, prior work in lifelong robot learning can be broadly classified into three distinct approaches. Replay methods (Brohan et al., 2023; Lopez-Paz & Ranzato, 2017; Chaudhry et al., 2019; Shin et al., 2017; Xie & Finn, 2022) retain previous data (e.g., vision-language web data, robot pretraining data, and data from seen tasks) and mix it with in-distribution data during training on new tasks. These methods leads to generalizable policies, but also necessitates huge storage space and computational overhead for retaining old samples. Regularization methods (Zenke et al., 2017; Kirkpatrick et al., 2017; Li & Hoiem, 2017) balance between new and old tasks by restricting the update of parameters. For example, EWC (Kirkpatrick et al., 2017) require an entire copy of frozen model from the previous task and entail substantial computation in estimating the importance of parameters. In addition, they usually struggle in long task sequence due to plasticity-stability dilemma.
In contrast, architectural methods (Rusu et al., 2016; Mallya & Lazebnik, 2018; Mallya et al., 2018; Ge et al., 2024) address catastrophic forgetting by explicitly learning task-specific parameters. The widely adopted parameter-efficient fine-tuning (PEFT) techniques (Ding et al., 2023), such as adapters (Houlsby et al., 2019; Chen et al., 2022), low-rank adaptation (LoRA) (Hu et al., 2021), and prompt tuning (Jia et al., 2022; Li & Liang, 2021; Lester et al., 2021), have proven highly effective for steering frozen foundation models by integrating small, learnable modules in single-task adaptation. In the lifelong learning setting, a practical solution, as exemplified by TAIL (Liu et al., 2024b), involves initializing and training task-specific adapters for each task, subsequently freezing them, and retrieving the appropriate adapter based on the oracle task index. Such method allows efficient adaptation with only a small amount of task-specific parameters, but also exhibits certain limitations. First, assuming test-time task identity is known and retrieving adapter in a hard-coded way could be impractical in real scenarios. Second, failing to effectively share knowledge across isolated adapters result in poor forward transfer. Some prior arts construct a shared adapter pool and enable automatic timestep-level retrieval by, for example, performing query-key matching to select the one with highest similarity from a pool of adapters (Schmied et al., 2023), or using multifaceted prototypes to represent the cluster center in the state space for each adapter (Lee et al., 2024), but usually suffer from inaccurate retrieval and suboptimal performance.
@@
In this paper, we introduce the Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) to address key challenges in lifelong robot learning. As illustrated in Figure 2, DMPEL features two core innovations: (1) it incrementally constructs a library of low-rank experts and employs a lightweight router to dynamically compose these experts into a unified policy based on the current context. This design allows the agent to flexibly adapt while leveraging all previously acquired parametric knowledge, thereby enhancing forward transfer; (2) it uses expert coefficient replay to regularize the router in accurately integrating experts for all previously encountered tasks, exploiting the modularity of fine-tuned parameters to prevent catastrophic forgetting. Notably, replaying low-dimensional embeddings and coefficients using the router is significantly more efficient than replaying demonstrations that require the entire policy in terms of storage and computation. Extensive evaluations on the lifelong manipulation benchmark LIBERO (Liu et al., 2024a) demonstrate that our method achieves superior forward transfer with reduced catastrophic forgetting compared to existing baselines, all while requiring minimal trainable parameters and storage.
Figure 2: Overview of the proposed method DMPEL.
@@ 2 Related Work
The traditional tabula rasa paradigm in robotics is widely criticized for its sample inefficiency, prompting researchers to explore the use of foundation models to improve downstream task transfer (Firoozi et al., 2023; Hu et al., 2023). One prominent direction involves transferring general visual representations: some studies directly employ foundation models from other domains as vision encoders (Parisi et al., 2022; Yuan et al., 2022), such as CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2024), while others pretrain visual representations on egocentric video datasets (Nair et al., 2023; Majumdar et al., 2024). Another emerging line of research focuses on pretraining large-scale policies on diverse robotic tasks to capture generalizable behavior priors, for example, RT-2 (Brohan et al., 2023), OpenVLA (Kim et al., 2025), and $\pi_{0}$ (Black et al., 2024). However, when adapting these models to downstream tasks, most approaches either rely on the frozen pretrained model (i.e., zero-shot transfer) or require costly full fine-tuning. Some prior work has explored parameter-efficient fine-tuning (PEFT) by integrating task-specific modules into the pretrained vision encoder (Sharma et al., 2023; Marza et al., 2024) or temporal transformer (Liang et al., 2022; Xu et al., 2023; Qiao et al., 2023). Despite these advancements, the majority of these arts focus on single-task adaptation, whereas our approach targets lifelong adaptation to a sequence of new tasks.
@@
Compared to full fine-tuning, PEFT optimizes only a small amount of inserted parameters while keeping the rest unchanged, which shares conceptual similarities with architectural methods in lifelong learning (Ding et al., 2023). Most studies on PEFT-based lifelong learning predominantly utilize prompt-based tuning (Jia et al., 2022; Lester et al., 2021) to adapt frozen pretrained model for tasks such as image classification, e.g., L2P (Wang et al., 2022c), DualPrompt (Wang et al., 2022b), CODA-Prompt (Smith et al., 2023), DAP (Jung et al., 2023), and text classification, e.g., EPI (Wang et al., 2023b), Progressive Prompts (Razdaibiedina et al., 2023). Current research primarily investigates strategies for dynamically selecting or generating instance-specific prompts (Zhou et al., 2024). Recognizing the representational limitations of prompt tuning, subsequent studies have extended to more powerful PEFT techniques, including LAE (Gao et al., 2023), HiDe-PET (Wang et al., 2025), and SKILL (Ge et al., 2024). Despite the success in incremental classification tasks, its application remains largely unexplored in robotics. A few pioneering works employ LoRA (Hu et al., 2021) to store task-specific knowledge, yet diverge in their retrieval mechanisms: TAIL (Liu et al., 2024b) relies on oracle task identifiers, L2M (Schmied et al., 2023) implements query-key matching between the current context and learnable keys, while IsCiL (Lee et al., 2024) establishes multifaceted prototypes through K-means clustering in the state space, with input states retrieving the closest LoRA parameters based on proximity. Our method incrementally constructs a parameter-efficient expert library and learns a lightweight router to dynamically retrieve and combine these experts.
@@
MoE approach involves training multiple specialized models for specific subtasks and has gained significant attention (Fedus et al., 2022; Muqeeth et al., 2024). A key method for combining these experts is parameter ensemble, supported by the linear mode connectivity phenomenon, which shows models converging to a single low-loss basin (Frankle et al., 2020). Task Arithmetic (Ilharco et al., 2023) introduces task vectors, representing the difference between fine-tuned and pretrained models, enhancing multi-task model performance when merged. Follow-up work (Yang et al., 2024; Tang et al., 2024) improves this by learning input-conditioned layer-wise fusion coefficients, increasing flexibility and performance. Model fusion techniques also apply to PEFT, where previously learned adapters are reused for efficient task adaptation (Wu et al., 2023; 2024; Wang et al., 2022a; Zhao et al., 2024), including LoRAHub (Huang et al., 2024) and AdapterFusion (Pfeiffer et al., 2021). In robotics, models like MELA (Yang et al., 2020), MoE-Loco (Huang et al., 2025), MORE (Zhao et al., 2025), and SDP (Wang et al., 2024b) create versatile, adaptive behaviors for multiple tasks by fusing expert sets. Our method aims to enhance forward transfer in lifelong robot learning by flexibly fusing previously learned low-rank experts for each sub-module in the policy according to the current context. Besides, we exploit the modularized design by replaying expert coefficient on the router to mitigate catastrophic forgetting.
@@ 3.1 Lifelong Robot Learning
Due to the significant challenge of sparse reward in robot manipulation, we consider a more practical lifelong imitation learning setting (Liu et al., 2024a; b), where a robot sequentially learns over a stream of tasks $\{\mathcal{T}^{1},\cdots,\mathcal{T}^{K}\}$ given an expert demonstration dataset $\mathcal{D}^{k}=\{\tau_{n}^{k}\}_{n=1}^{N}$ for each task $\mathcal{T}^{k}=(d_{0}^{k},l^{k})$, $1\leq k\leq K$. Each expert trajectory $\tau_{n}^{k}$ can be represented as $(o_{0},a_{0},\cdots,o_{\mathcal{H}})$, in which observation $o_{t}$ at each step $t$ includes RGB images and proprioceptive states. Following common pratice in partially observable MDPs, we approximate the current state $s_{t}$ by stacking prior observations, i.e., $s_{t}=o_{\leq t}=(o_{0},\cdots,o_{t})$. We perform behavior cloning (Bain & Sammut, 1995) to learn a stochastic policy $\pi_{\theta}$ by minimizing the negative log-likelihood loss $L$:
@@ 3.2 Evaluation Metrics
We use three metrics to evaluate the performance of lifelong robot learning (Liu et al., 2024a; Rodríguez et al., 2018): forward transfer (FWT), negative backward transfer (NBT), area under the success rate curve (AUC). Formally, we evaluate the policy on current task $\mathcal{T}^{k}$ when reaching a checkpoint $c\in\mathcal{C}$ (i.e., every multiple epochs of training) and we denote the success rate during evaluation as $s_{k,k,c}$. When the training on current task is completed, we find the earliest checkpoint $c_{k}^{*}$ that achieves the best performance and mark it as the final success rate $s_{k,k}$, i.e., $s_{k,k}=s_{k,k,c_{k}^{*}}=\max_{c\in\mathcal{C}}s_{k,k,c}$. We only keep the best checkpoint for future training and evaluation as if the agent stops learning after $c_{k}^{*}$, i.e., for all $c\geq c_{k}^{*}$, $s_{k,k,c}=s_{k,k}$. This checkpoint $c_{k}^{*}$ will be further evaluated on previous tasks $\mathcal{T}^{j},j\leq k$, where the success rate is denoted as $s_{k,j}$. Intuitively, higher FWT indicates faster adaptation to new tasks, lower NBT indicates less catastrophic forgetting on old tasks, and higher AUC demonstrates an overall better performance across tasks and a good balance between FWT and NBT. These metrics can be formally defined as:
@@ 4.1 Base Policy Architecture
We use a policy similar to the one used in the LIBERO benchmark (Liu et al., 2024a; b) and focus on developing better lifelong learning algorithms. The policy consists of vision, text, and state encoders, an input modality fusion module, a temporal transformer, and an action head.
@@
The policy input is the historical observations $(o_{t-T},\cdots,o_{t})$ and the language instruction $l=\{l_{i}\}_{i=1}^{L}$, where $T$ is the context length and $L$ is the sentence length. The observation at each step $o_{t}=(I^{1}_{t},\cdots,I^{N_{c}}_{t},p_{t})$ includes RGB images $I^{n_{c}}_{t},1\leq n_{c}\leq N_{c}$ from $N_{c}$ different cameras and low-dimensional proprioceptive states $p_{t}$. In experiments, we use images from two cameras, the agent-view image and the eye-in-hand image. We employ the pretrained CLIP ViT-B/16 model (Radford et al., 2021) to serve as the vision encoder $\mathcal{E}_{I}$ and the text encoder $\mathcal{E}_{L}$. For the joint states and gripper states, we learn separate linear projection layers $\mathcal{E}_{P}$ to project the low-dimensional states to proprioceptive embeddings. In summary, at each timestep, we obtain image embeddings $f^{v}_{t}$, text embeddings $f^{l}_{t}$, and proprioceptive embeddings $f^{s}_{t}$.
@@
We utilize a feature-wise linear modulation (FiLM) (Perez et al., 2018) to fuse language embeddings with the image and state embeddings. Generally, given the original feature $x$, the conditional input $z$, and the fusion network $f_{\text{FiLM}}$, we can obtain the modulated feature $x^{\prime}=\gamma\odot x+\beta$ where $\gamma,\beta=f_{\text{FiLM}}(z)$. In our case, the conditional input $z$ refers to the language embedding $f^{l}_{t}$, $\gamma$ and $\beta$ are scaling and shifting vectors having the same size as $x$, and $\odot$ refers to element-wise multiplication.
@@
The causal temporal transformer backbone $\mathcal{D}_{T}$ processes a sequence of multi-modal embeddings to produce a latent vector $g_{t}$ at each decision-making timestep. We employ a Gaussian Mixture Model (GMM)-based output head $\mathcal{D}_{H}$ to compute the multi-modal distribution of manipulation actions (Liu et al., 2024b; a). During training, we optimize the negative log-likelihood loss between the action distribution and the ground truth action. At inference time, the robot executes the policy by using the mean of the Gaussian distribution with the highest density for end effectors.